OpenStudy (anonymous):

Help me please, Does the following equation represent growth or decay? y = 2(4)-x

OpenStudy (anonymous):

i dont think so growth/decay equations contain exponents

OpenStudy (anonymous):

y = 2(4)^-x <---- sorry that's it

OpenStudy (anonymous):

Given y=2(4)^-x, rearrange it to the follow: \[y= 2 (\frac{1}{4})^{x}\] Looking at the equation above, when x increases y will be come smaller and smaller, because of the (1/4) component. When x = 1, y= 2(1/4) WHen x = 2, y= 2(1/16) When x = 3, y= 2(1/ 64). Since y decreases exponentially while x increases it is a decay function.

OpenStudy (anonymous):

Thank you